OpenStudy (anonymous):

Which of the following is a true statement? Choose one answer. A. All square roots are irrational. B. All decimals are irrational. C. All non-perfect squares have irrational square roots. D. All irrational numbers can be written as ratios of two integers.

OpenStudy (anonymous):

C is true

OpenStudy (anonymous):

A) \[\sqrt{4}\] is rational, so A is false B) 0.1 can be written as \[\frac{ 1 }{ 10 }\]which is rational, so B is false D) Pi cannot be written as a fraction, so D is false, which leaves C C) \[\sqrt{3}\] has no rational square roots, so C is true